Page layout in L^{A}TEX

Piet van Oostrum^{∗} Dept. of Computer Science Utrecht University

June 26, 1996

Abstract This article describes how to customize the page layout of your LaTeX documents, i.e how

to change page margings and sizes, headers and footers, and the proper placement of figures and tables (collectively called floats) on the page.

Originally this was the documentation of the fancyheadings package. It did contain also other info, e.g. advanced use of marks. It has now been upgraded to include more, e.g. the handling of floats. The fancyheadings documentation has been upgraded to conform to version 2 of this package^{1}. For reasons of compatibility with certain operating systems, the name of the package has been changed to fancyhdr.

Although this paper uses L^{A}TEX 2ε commands, most of the techniques can be used with older L^{A}TEXversions with appropriate changes.

Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2 Page headers and footers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3 What is fancyhdr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4 Simple use of fancyhdr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 5 A simple example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 6 An example of two-sided printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 7 Redefining plain style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 8 The default layout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 9 The scoop on L^{A}TEX’s marks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 10 Dictionary style headers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 11 Fancy layouts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 12 Two book examples . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 13 Special page layout for float pages . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 14 Those blank pages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 15 N of M style page numbers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 16 Chapter or section related page numbers . . . . . . . . . . . . . . . . . . . . . . . . 16 17 When to change the headers and footers? . . . . . . . . . . . . . . . . . . . . . . . 16

∗A considerable part of this article was written by George Grätzer (University of Manitoba) in Notices Amer.

Math. Soc. Thanks, George!

1this version is due to be released Real Soon Now

18 Headers and footers induced by the text . . . . . . . . . . . . . . . . . . . . . . . . 17

19 A movie . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

20 Thumb-indexes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

21 Float placement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20

22 Multipage Floats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

22.1 Tables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23

22.2 Figures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24

23 Contact information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

1 Introduction

A page in a L^{A}TEX document is built from various elements as shown in figure 1. The body contains the main text of the document together with the so called floats (tables and figures).

The pages are constructed by L^{A}TEX’s output routine, which is quite complicated and should therefore not be modified. Some of the packages described in this paper contains small modifications to the output routine to accomplish things that cannot be done in another way. You should use these packages to get the desired result rather than fiddling with the output routine yourself.

There a a number of things that you must be aware of:

1. The margins on the left are not called \leftmargin, but \evensidemargin (on evennumbered pages) and \oddsidemargin (on odd-numbered pages). In one-sided documents \oddsidemargin is used for either. \leftmargin is also a valid L^{A}TEX parameter but it has a different use (namely the indentation of lists).

2. Most of the parameters should not be changed in the middle of a document. Some changes might work at a pagebreak. If you want to change the height of a single page, you can use

     the \enlargethispage command.

The margin notes area contains small pieces of information created by the \marginpar command. On twosided documents the margin notes appear on the left and right alternatively. The margin notes are not on fixed places with respect to the paper but at approximately the same height as the paragraph in which they appear. Due to the algorithm used to decide the placement of margin notes, in a twosided document unfortunately they may appear on the wrong side if they are close to a page break. If you want to put information on fixed places in the margins you may use the technique described in sections 19 and 20.

The first part of this paper describes how to change the header and footer areas. The last part describes how to get your floats at the desired place.

2 Page headers and footers

The page headers and footers in L^{A}TEX are defined by the \pagestyle and \pagenumbering commands. \pagestyle defines the general contents of the headers and footers (e.g. where the page number will be printed), while \pagenumbering defines the format of the page number. L^{A}TEX has four standard pagestyles:

[figure: Header; Body; Footer; Margin Notes; 8; 7; 1; 3; 10; 9; 11; 2; 4; 5; 6]

1   one inch + \hoffset        2   one inch + \voffset
3   \oddsidemargin = 18pt      4   \topmargin = 0pt
5   \headheight = 12pt         6   \headsep = 25pt
7   \textheight = 646pt        8   \textwidth = 424pt
9   \marginparsep = 11pt      10   \marginparwidth = 54pt
11   \footskip = 30pt               \marginparpush = 5pt (not shown)
    \hoffset = 0pt                 \voffset = 0pt
    \paperwidth = 597pt            \paperheight = 845pt

Figure 1: Page elements. The values shown are those in effect in the current document, not the defaults.

empty  no headers or footers
plain  no header, footer contains page number centered

headings no footer, header contains name of chapter/section and/or sub-

section and page number

myheadings no footer, header contains page number and user supplied information

Although these are useful styles, they are quite limited. Additional page styles can be defined by defining commands of the form \ps@xxx. This command is executed when a \pagestyle{xxx} is given in the document. The \ps@xxx command should define the following commands for the contents of the headers and footers: \@oddhead header on odd numbered pages in two-sided documents (on all

pages in one-sided)

\@evenhead header on even numbered pages in two-sided documents

\@oddfoot footer on odd numbered pages in two-sided documents (on all

pages in one-sided)

\@evenfoot footer on even numbered pages in two-sided documents

These are not user commands, but rather “variables” that are used by L^{A}TEX’s output routine. As the command names contain the character ’@’, they should be defined in a package file, or otherwise be sandwiched between the commands \makeatletter and \makeatother.

The \pagenumbering command defines the layout of the page number. It has a parameter from the following list:

arabic arabic numerals
roman lower case roman numerals
Roman upper case roman numerals
alph  lower case letter
Alph  upper case letter

The \pagenumbering{xxx} defines the command \thepage to be the expansion of the page number in the given notation xxx. The pagestyle command then would include \thepage in the appropriate place. Additionally the \pagenumbering command resets the page number to 1. The \pagestyle and \pagenumbering apply to the page that is being constructed, so they should be used at a location where it is clear to what page they apply (see section 17).

3 What is fancyhdr

The fancyhdr macro package allows you to customize in L^{A}TEX your page headers and footers in an easy way. You can define:

• three-part headers and footers

• decorative lines in headers and footers

• headers and footers wider than the width of the text

• multi-line headers and footers

• separate headers and footers for even and odd pages

• different headers and footers for chapter pages

• different headers and footer on pages with floats

Of course, you also have complete control over fonts, uppercase and lowercase displays, etc.

4 Simple use of fancyhdr

To use this package in a L^{A}TEX 2ε document, place the file fancyhdr.sty in a directory/folder where TEX can find it (normally in the input directory/folder), and include in the preamble of your document after

\documentclass{...}

the commands^{2}:

\usepackage{fancyhdr}
\pagestyle{fancy}

We can visualize the page layout we can create with fancyhdr as follows:

LeftHeader  CenteredHeader  RightHeader

page body

LeftFooter  CenteredFooter  RightFooter

The LeftHeader and LeftFooter are left justified; the CenteredHeader and CenteredFooter are centered; the RightHeader and RightFooter are right justified.

We define each of the six “fields” and the two decorative lines separately.

5 A simple example

K. Grant is writing a report to Dean A. Smith, on “The performance of new graduates” with the following page layout:

The performance of new graduates

page body

From: K. Grant  To: Dean A. Smith  3

where “3” is the page number. The title: “The performance of new graduates” is bold.

This is accomplished by these commands following \pagestyle{fancy} ^{3}:

\lhead{}
\chead{}
\rhead{\bfseries The performance of new graduates}
\lfoot{From: K. Grant}
\cfoot{To: Dean A. Smith}

2For LaTeX2.09 you should specify [fancyhdr] in the \documentstyle instead of the \usepackage command.

3Note that version 1 of fancyheadings used the \setlength command to change the \...rulewidth parameters.

\rfoot{\thepage}
\renewcommand{\headrulewidth}{0.4pt}
\renewcommand{\footrulewidth}{0.4pt}

(The \thepage macro displays the current page number. \bfseries is the L^{A}TEX 2ε’s way of selecting bold face.)

This is now fine, except that the first page does not need all these headers and footers. To eliminate all but the centered page number, issue the command

\thispagestyle{plain}
after the \begin{document} and the \maketitle commands.

Alternatively, issue

\thispagestyle{empty}

if you do not want any headers or footers.

In fact the standard L^{A}TEX classes have the command \maketitle defined in such a way that a \thispagestyle{plain} is automatically issued. So if you do want the fancy layout on a page

containing \maketitle you must issue a \thispagestyle{fancy} after the \maketitle.

6 An example of two-sided printing

Some document classes, such as book.cls, print two-sided by default: the even pages and the odd pages have different layouts; other document classes use the twoside option to print two-sided.

Now let us print the report two-sided. Let the above page layout be used for the odd (right-side) pages, and the following for the even (left-side) pages:

The performance of new graduates

page body

4  From: K. Grant  To: Dean A. Smith

where “4” is the page number.

Here are the commands:

\fancyhead{} % clear all fields
\fancyhead[RO,LE]{\bfseries The performance of new graduates}
\fancyfoot[LE,RO]{\thepage}
\fancyfoot[LO,CE]{From: K. Grant}
\fancyfoot[CO,RE]{To: Dean A. Smith}
\renewcommand{\headrulewidth}{0.4pt}
\renewcommand{\footrulewidth}{0.4pt}

We use the more general commands \fancyhead and \fancyfoot. These have an additional parameter between square brackets that specifies for which pages and/or parts of the header/footer they apply. The first \fancyhead command omits this parameter, and thus applies to all header

E Even page
O Odd page
L Left field
C Center field
R Right field
H Header
F Footer

Figure 2: Selectors

fields. In general this is only useful to get rid of the defaults or a previous definition, as is done here. The selectors that can be used between the square brackets are given in figure 2. Selectors can be combined so \fancyhead[LE,RO]{text} will define the field for both the left header on even pages and the right header on odd pages. If you don’t give an E or O the definition applies to both. Similar for LRC. So the use of \lhead in the previous section is just an abbreviation for \fancyhead[L]. The selectors may be given as uppercase or lowercase letters.

There is also a more general command \fancyhf that you can use to combine the specifications for headers and footers. This allows additional selectors H (header) and F (footer). In fact \fancyhead and \fancyfoot are just \fancyhf with H and F prespecified.

Again, you may use \thispagestyle{plain} for a simple page layout for page 1.

7 Redefining plain style

Some L^{A}TEX commands, like \chapter, use the \thispagestyle command to automatically switch to the plain page style, thus ignoring the page style currently in effect. To customize even such pages you must redefine the plain pagestyle. As we indicated before you could do this by defining the \ps@plain command, but fancyhdr gives you an easier way with the \fancypagestyle command. This command can be used to redefine existing pagestyles (like plain) or to define new ones, e.g. if part of your document is to use a different pagestyle. This command has two parameters: one is the name of the pagestyle to be defined, the second consists of commands that change the headers and/or footers, i.e. fancyhead etc. Also allowed are changes to \headrulewidth and \footrulewidth. As an example, let us redefine the plain style for the report in Section 6 by making the page number bold.

\fancypagestyle{plain}{%
\fancyhf{} % clear all header and footer fields
\fancyfoot[C]{\bfseries \thepage} % except the center
\renewcommand{\headrulewidth}{0pt}
\renewcommand{\footrulewidth}{0pt}}

8 The default layout

Let us use the book.cls documentclass and the default settings for fancyhdr; so we only issue the commands

\usepackage{fancyhdr}
\pagestyle{fancy}

and let fancyhdr take care of everything. On the pages where new chapters start, we get a centered page number in the footer; there is no header, and there are no decorative lines.

On an even page, we get the layout:

1.2 EVALUATION CHAPTER 1. INTRODUCTION

page body

4

On an odd page, we get the layout:

CHAPTER 1. INTRODUCTION 1.2 EVALUATION

page body

3

where the header text is slanted uppercase.

This default layout is produced by the following commands:

\fancyhead[LE,RO]{\slshape \rightmark}
\fancyhead[LO,RE]{\slshape \leftmark}
\fancyfoot[C]{\thepage}

The following settings are used for the decorative lines:

 \headrulewidth  0.4pt
 \footrulewidth  0 pt

The header text is turned into all uppercase in book.cls.

9 The scoop on L^{A}TEX’s marks

Usually, for documents of class book and report, you may want to use chapter and section information in the headings (chapter only for one-sided printing), and for documents of class article, section and subsection information (section only for one-sided printing). L^{A}TEX uses a marker mechanism to remember the chapter and section (section and subsection) information for a page; this is discussed in detail in the L^{A}TEX Companion, Section 4.3.1.

There are two ways you can use and change the higher- and lower-level sectioning information available to you. The macros: \leftmark (higher-level) and \rightmark (lower-level) contain the information processed by L^{A}TEX, and you can use them directly as shown in Section 8.

The \leftmark contains the Left argument of the Last \markboth on the page, the \rightmark contains the Right argument of the fiRst \markboth or the only argument of the fiRst \markright on the page. If no marks are present on a page they are “inherited” from the previous page.

You can influence how chapter, section, and subsection information (only two of them!) is displayed

by redefining the \chaptermark, \sectionmark, and \subsectionmark commands^{4}. You must

put the redefinition after the first call of \pagestyle{fancy} as this sets up the defaults.

4There are similar commands for paragraph and subparagraph but they are seldom used.

Code:

\renewcommand{\chaptermark}[1]{%
 \markboth{\chaptername
 \ \thechapter.\ #1}{}}
\renewcommand{\chaptermark}[1]{%
 \markboth{\MakeUppercase{%
 \chaptername}\ \thechapter.%
 \ #1}{}}
\renewcommand{\chaptermark}[1]{%
 \markboth{\MakeUppercase{%
 \chaptername\ \thechapter.%
 \ #1}}{}}
\renewcommand{\chaptermark}[1]{%
 \markboth{#1}{}}
\renewcommand{\chaptermark}[1]{%
 \markboth{\thechapter.\ #1}{}}
\renewcommand{\chaptermark}[1]{%
 \markboth{\thechapter.%
 \ \chaptername.\ #1}{}}

Prints:

Chapter 2. Do it now

CHAPTER 2. Do it now

CHAPTER 2. DO IT NOW

Do it now

2. Do it now

2. Chapter. Do it now

Figure 3: Marker variants

Let us illustrate this with chapter info. It is made up of three parts:

• the number (say, 2), displayed by the macro \thechapter

• the name (in English, Chapter), displayed by the macro \chaptername

• the title, contained in the argument of \chaptermark.

Figure 3 shows some variants for “Chapter 2. Do it now” (the last example is appropriate in some non-English languages). The % signs at the end of the lines are to prevent unwanted space. Normally you would continue the lines and remove these % signs^{5}.

For the lower-level sectioning information, do the same with \markright.

So if “Section 2.2. First steps” is the current section, then

\renewcommand{\sectionmark}[1]{\markright{\thesection.\ #1}}

will give “2.2. First steps”

Redefining the \chaptermark and \sectionmark commands may not eliminate all uppercaseness. E.g. the bibliography will have a title of bibliography in the header, as the \MakeUppercase is explicitly given in the definition of \thebibliography. Similar for index etc. If you don’t want to redefine these commands, you can use the \nouppercase command that fancyhdr makes available in the header and footer fields. Note that this may screw other things, like uppercase roman numerals in your headers, so it should be used with care. Essentially this command typesets its argument in an environment where \MakeUppercase and \uppercase are changed into do-nothing operations.

5the \MakeUppercase command is used in L^{A}TEX 2ε to generate uppercase text, while in L^{A}TEX 2.09 \uppercase is used. The difference is that \MakeUppercase also deals with non-ASCII letters. Fancyhdr defines \MakeUppercase to be an alias for \uppercase if it isn’t defined.

\lhead{\nouppercase{\rightmark}}
\rhead{\nouppercase{\leftmark}}

It should be noted that the L^{A}TEX marking mechanism works fine with chapters (which always start on a new page) and sections (which are reasonably long). It does not work quite as well with short sections and subsections. This is a problem with L^{A}TEX, not with fancyhdr.

As an example let’s take a page layout where the leftmarks are generated by the sections and the rightmarks by the subsections (as is default in the article class). Take a page with some short sections, e.g.

Section 1. subsection 1.1 subsection 1.2 Section 2.

As the leftmark contains the last mark of the page it will be “Section 2.”, and the rightmark will be “subsection 1.1” as it will be the first mark of the page. So the page header info will combine section 2 with subsection 1.1 which isn’t very nice. The best you can do in these cases is use only the \rightmarks and redefine \sectionmark accordingly. A L^{A}TEX command \firstleftmark would also be a nice addition (see the extramarks package in section 18).

Another problem with the marks in the standard L^{A}TEX classes is that the higher level sectioning commands (e.g. \chapter) call \markboth with an empty right argument. This means that on the first page of a chapter (or a section in article style) the \rightmark will be empty. If this is a problem you must manually insert extra \markright commands or redefine the \chaptermark (\sectionmark) commands to issue a \markboth command with two decent parameters.

As a final remark you should also note that the * forms of the \chapter etc. commands do not call the mark commands. So if you want your preface to set the header info but not be numbered nor be put in the table of contents, you must issue the \markboth command yourself, e.g.

\chapter*{Preface\markboth{Preface}{}}

Entering the \markboth command inside the \chapter* insures that the mark will not be separated from the title by a page break. Of course with \chapter* this wouldn’t be a problem if you put the mark command after the chapter title, as the \chapter* command starts at a new page. However with a \section* it could be dangerous to say:

\section*{Preface}
\markboth{Preface}{}

as a page break may occur between the two commands.

10 Dictionary style headers

Dictionaries and concordances usually have a header containing the first word defined on the page or both the first and the last words. This can easily be accomplished with fancyhdr and L^{A}TEX’s mark mechanism. Of course if you use the marks for dictionary style headers, you cannot use them for chapter and section information, so if there are also chapters and sections present, you must redefine the \chaptermark and \sectionmark to make them harmless:

% fixmarks.sty:
% Patch LaTeX's output routine to handle marks correctly with two columns.
% Joe Pallas <pallas@edu.stanford.neon>
% Corrected by Piet van Oostrum <piet@cs.ruu.nl> on Feb 5, 1993, Oct 5, 1994
\def\@outputdblcol{\if@firstcolumn \global\@firstcolumnfalse
% Remember the marks from the first column
    \global\setbox\@leftcolumn\copy\@outputbox
    \splitmaxdepth=\maxdimen \cbaddness=10000
    \setbox\@outputbox\vsplit\@outputbox to\maxdimen
    \xdef\@firstcoltopmark{\topmark}%
    \xdef\@firstcolfirstmark{\splitfirstmark}%
    \ifx\@firstcolfirstmark\empty\global\let\@setmarks\relax\else
      \gdef\@setmarks{\let\firstmark\@firstcolfirstmark
        \let\topmark\@firstcoltopmark}%
    \fi
% End of change
  \else \global\@firstcolumntrue
    \setbox\@outputbox\vbox{\hbox to\textwidth{\hbox to\columnwidth
      {\box\@leftcolumn \hss}\hfil \vrule width\columnseprule\hfil
       \hbox to\columnwidth{\box\@outputbox \hss}}}\@combinedblfloats
% Override current first and top with those of first column if necessary
       \@setmarks
% End of change
       \@outputpage \begingroup \@dblfloatplacement \@startdblcolumn
       \@whilesw\if@fcolmade \fi{\@outputpage\@startdblcolumn}\endgroup
    \fi}

Figure 4: Fix marks in twocolumn style

\renewcommand{\chaptermark}[1]{}
\renewcommand{\sectionmark}[1]{}

Now you do a \markboth{#1}{#1} for each dictionary or concordance entry #1 and use \rightmark for the first entry defined on the page and \leftmark for the last one.

If you want to use a header entry of the form firstword–lastword it would be nice if this would be reduced to just the form firstword if both are the same. This could happen if there is just one entry on the page. In this case a test must be made to check if the marks are the same. However, TEX’s marks are strange beasts, which cannot be compared out of the box with the plain TEX \if commands. Fortunately the ifthen package works well:

\newcommand{\mymarks}{
  \ifthenelse{\equal{\leftmark}{\rightmark}}
    {\rightmark} % if equal
    {\rightmark--\leftmark}} % if not equal
\fancyhead[LE,RO]{\mymarks}
\fancyhead[LO,RE]{\thepage}

Dictionaries are often done with two columns. Unfortunately there is a bug in L^{A}TEX’s twocolumn option which causes some marks to be lost. If you use the package in figure 4, this will be solved^{6}^{7}.

6This is a simplification of the actual code. Get the real package from the CTAN sites.

7The multicol package uses a similar technique.

11 Fancy layouts

You can make a multi-line field with the \\ command. It is also possible to put extra space in a field with the \vspace command. Note that if you do this you will probably have to increase the height of the header (\headheight) and/or of the footer (\footskip), otherwise you may get error messages “Overfull \vbox … has occurred while \output is active”^{8}. See Section 4.1 of the L^{A}TEX Companion for detail.

For instance, the following code will place the section title and the subsection title of an article in two lines in the upper right hand corner:

\documentclass{article}
\usepackage{fancyhdr}
\pagestyle{fancy}
\addtolength{\headheight}{\baselineskip}
\renewcommand{\sectionmark}[1]{\markboth{#1}{}}
\renewcommand{\subsectionmark}[1]{\markright{#1}}
\rhead{\leftmark\\\rightmark}

You can also customize the decorative lines. You can make the decorative line in the header quite thick with

\renewcommand{\headrulewidth}{0.6pt}

or you can make the decorative line in the footer disappear with

\renewcommand{\footrulewidth}{0pt}

The decorative lines, themselves, are defined in the two macros \headrule and \footrule. For instance, if you want a dotted line rather than a solid line in the header, redefine the command

\headrule:
\renewcommand{\headrule}{\vbox to 0pt{\hbox
    to\headwidth{\dotfill}\vss}}

There is one additional parameter that you can set: \footruleskip. It defines the distance between the decorative line in the footer and the top of the footer text line. By default it is set to 30% of the normal line distance. You may want to adjust it if you use unusally large or small fonts in the footer. Change it with \renewcommand.

12 Two book examples

The following definitions give an approximation of the style used in L. Lamport’s L^{A}TEX book.

Lamport’s header overhangs the outside margin. This is done as follows.

The width of headers and footers is \headwidth, which by default equals the width of the text: \textwidth. You can make the width wider (or narrower) by redefining \headwidth with the \setlength and \addtolength commands. To overhang the outside margin where the marginal

notes are printed, add both \marginparsep and \marginparwidth to \headwidth with the

commands:

8If you use 11pt or 12pt you will probably also have to do this, because L^{A}TEX’s defaults are quite small

\addtolength{\headwidth}{\marginparsep}
\addtolength{\headwidth}{\marginparwidth}

You must issue these commands after the first \pagestyle{fancy} command as this will establish the default for \headwidth^{9}.

And now a complete definition of Lamport’s book style:

\documentclass{book}
\usepackage{fancyhdr}
\pagestyle{fancy}
\addtolength{\headwidth}{\marginparsep}
\addtolength{\headwidth}{\marginparwidth}
\renewcommand{\chaptermark}[1]{\markboth{#1}{}}
\renewcommand{\sectionmark}[1]{\markright{\thesection\ #1}}
\fancyhf{}
\fancyhead[LE,RO]{\bfseries\thepage}
\fancyhead[LO]{\bfseries\rightmark}
\fancyhead[RE]{\bfseries\leftmark}
\fancypagestyle{plain}{%
   \fancyhead{} % get rid of headers
   \renewcommand{\headrulewidth}{0pt} % and the line
}

Notice that the \chaptermark and \sectionmark commands have been redefined to eliminate the chapter numbers and the uppercaseness.

For the second example, we take the A_{M}S-L^{A}TEX book.

Chapter pages have no headers or footers. So we declare

\thispagestyle{empty}

for every chapter page, and we do not need to redefine plain.

Chapter and section titles appear in the form: 2. DO IT NOW, so we have to redefine \chaptermark and \sectionmark as follows (see Section 9):

\renewcommand{\chaptermark}[1]%
   {\markboth{\MakeUppercase{\thechapter.\ #1}}{}}
\renewcommand{\sectionmark}[1]%
   {\markright{\MakeUppercase{\thesection.\ #1}}}

In an even-header, the page number is printed as the LeftHeader and the chapter info as the RightHeader; in an odd-header, the section info is printed as the LeftHeader and the page number as the RightHeader. The CenteredHeaders are empty. There are no footers.

There is a decorative line in the header. It is 0.5pt wide, so we need the commands:

\renewcommand{\headrulewidth}{0.5pt}
\renewcommand{\footrulewidth}{0pt}

The font used in the headers is 9 pt bold Helvetica. The PSNFSS system by Sebastian Rahtz uses the short (Karl Berry) name phv for Helvetica, so this font is selected with the commands:

9In version 2 of fancyhdr it is usually safe to do it before, but it is still advised to do after.

\fontfamily{phv}\fontseries{b}\fontsize{9}{11}\selectfont

(See Sections 7.6.1 and 11.9.1 of the L^{A}TEX Companion.) Let us define a shorthand for this:

\newcommand{\helv}{%
   \fontfamily{phv}\fontseries{b}\fontsize{9}{11}\selectfont}

Now we are ready for the page layout:

\documentclass{book}
\usepackage{fancyhdr}
\pagestyle{fancy}
\renewcommand{\chaptermark}[1]%
   {\markboth{\MakeUppercase{\thechapter.\ #1}}{}}
\renewcommand{\sectionmark}[1]%
   {\markright{\MakeUppercase{\thesection.\ #1}}}
\renewcommand{\headrulewidth}{0.5pt}
\renewcommand{\footrulewidth}{0pt}
\newcommand{\helv}{%
   \fontfamily{phv}\fontseries{b}\fontsize{9}{11}\selectfont}
\fancyhf{}
\fancyhead[LE,RO]{\helv \thepage}
\fancyhead[LO]{\helv \rightmark}
\fancyhead[RE]{\helv \leftmark}

13 Special page layout for float pages

Some people want to have a special layout for float pages (pages only containing floats). As these pages are generated autonomically by L^{A}TEX, the user doesn’t have any control over them. There is no \thispagestyle for float pages and any change of the page style will at least also affect the page before the float page. With fancyhdr, however, you can specify in each of the header- or footer fields

\iffloatpage{value for float page}{value for other pages}

You can even use this to get rid of the decorative line on float pages only by defining:

\renewcommand{\headrulewidth}{\iffloatpage{0pt}{0.4pt}}

Sometimes you may want to change the layout also for pages that contain a float on the top of the page or a float on the bottom of the page.

fancyhdr gives you the commands \iftopfloat and \ifbotfloat similar to \iffloatpage.

Note: Marks in floats will not be visible in L^{A}TEX’s output routine, so it is not useful to put marks in floats. So there is currently no way to let a float (e.g. a figure caption) influence the page header or footer.

14 Those blank pages

In the book class when the openany option is not given or in the report class when the openright option is given, chapters start at odd-numbered pages, half of the time causing a blank page to be inserted. Some people prefer this page to be completely empty, i.e. without headers and footers. This cannot be done with \thispagestyle as this command would have to be issued on the previous page. There is, however, no magic necessary to get this done:

\clearpage{\pagestyle{empty}\cleardoublepage}

As the \pagestyle{empty} is enclosed in a group it only affects the page that may be generated by the \cleardoublepage. You can of course put the above in a private command. If you want to have this done automatically at each chapter start or when you want some other text on the page then you must redefine the \cleardoublepage command.

\makeatletter
\def\cleardoublepage{\clearpage\if@twoside \ifodd\c@page\else
  \hbox{}
  \vspace*{\fill}
  \begin{center}
    This page intentionally contains only this sentence.
  \end{center}
  \vspace{\fill}
  \thispagestyle{empty}
  \newpage
  \if@twocolumn\hbox{}\newpage\fi\fi\fi}
\makeatother

15 N of M style page numbers

Some document writers prefer the pages to be numbered as n of m where m is the number of pages in the document. There is a package nofm.sty available, but some versions of it are defective, and most don’t work with fancyhdr because they take over the complete page layout. For L^{A}TEX 2ε there is a package lastpage available which you can use with fancyhdr as follows:

\usepackage{lastpage}
...
\cfoot{\thepage\ of \pageref{LastPage}}

If you are still using L^{A}TEX2.09 and you are not able to switch to L^{A}TEX 2ε you can use the L^{A}TEX2.09 compatible lastpage209.sty. which is defined as follows:

\let\origenddocument=\enddocument
\def\enddocument{\clearpage\if@filesw
  {\addtocounter{page}{-1} \immediate\write\@mainaux
    {\string\newlabel{LastPage}{{}{\thepage}}}}\origenddocument}

The value of the LastPage label can be used to make different headers or footers on the last page of a document. E.g. if you want the footer of every odd page, except if it is the last one, to contain the text “please turn over”, this can be done as follows^{10}:

\usepackage{lastpage}
\usepackage{ifthen}
...
\rfoot{\ifthenelse{\isodd{\value{page}} \and \not
    \value{page}=\pageref{LastPage}{please turn over}{}}

10This requires a reasonably recent version of the ifthen package.

16 Chapter or section related page numbers

In technical documentation very often page numbers are used of the form 2-10 where the first number is the chapter number and the second is the pagenumber relative to the chapter. Sometimes section is used rather than chapter. The package chappg can be used to get this format. If you want to change the layout e.g. use a dot rather than a dash, or a section rather than a chapter, you should make a private copy and edit it.

This package redefines \thepage as \arabic{chapter}-\arabic{page}.  Unfortunate-

ly this gives numbers rather than letters for appendices. A better definition would be \thechapter-\arabic{page} but you can give this definition yourself after the \usepackage{chappg} command. What the package also does is reset the page number to 1 at the beginning of each chapter.

There is a fundamental difference between the page numbering of the style “m of n” as described in the previous section and the current one. The m of n style is only used in the page header or footer, but not in the table of contents, index, or references like “See page xx”. Therefore it does not change the command \thepage. The page numbering style “2-10”, however should be used in all references to the page number, therefore it must be done by redefining \thepage.

17 When to change the headers and footers?

Sometimes you want to change the header or footer layout in the course of a document. Some of these changes can be accomplished by using the mark mechanism as may be seen in section 9 and 18. However, sometimes we want a more drastic change, e.g to change the page numbering from roman to arabic (with \pagenumbering), to change one of the fancyhdr fields or to change to another page style. Sometimes you may be surprised to find the change to occur too early. In general the above mentioned changes take effect immediately, i.e. on the page that is currently being built. If you want the change to take effect at the next page you must make sure that the current page is finished. In most cases this can be done by issuing a \clearpage command before any of the above mentioned changes. If this is not possible you can use the afterpage package with:

\afterpage{\lhead{new value}} or \afterpage{\pagenumbering{roman}. You cannot use

\afterpage to change the \pagestyle as the commands issued by \afterpage are local in a group, and the \pagestyle command makes only local changes. The \pagenumbering and the fancyhdr commands make global changes so they will work, as will the \thispagestyle command.

It should be noted that although the fancyhdr commands like \fancyhead take effect immediately, this does not mean that any “variables” used in these commands get the value they have at the place where these commands are given. E.g. if \fancyfoot[C]{\thepage} is given the page number that will be inserted in the footer is not the page number of the page where this command is given, but rather the page number of the actual page where the footer is constructed. Of course for the page number this is what you expect, but it is also true for other commands.

So if you have a book where each chapter is written by a different author and you want the name of the author in the lower left-hand corner you can use the following commands:

\newcommand{\TheAuthor}{}
\newcommand{\Author}[1]{\renewcommand{\TheAuthor}{#1}}
\lfoot{\TheAuthor}

and start each chapter with the command \Author{Real Name}. If however, the author name would be changed before a page is completed the wrong author could come in the footer. This would be the case if you gave the above command before the \chapter command rather than after it. Another source of problems is the fact that TEX’s output routine processes commands ahead, so it may already have processed some commands that produce text that will appear on the next page. See the next section for an example.

18 Headers and footers induced by the text

We have seen how we can use L^{A}TEX’s marks to get information from the document contents to the headers and footers. The marks mechanism is the only reliable mechanism that you can use to get changing information to the headers or footers. This is because L^{A}TEX may be processing your document ahead before deciding to break the page.

Sometimes the two marks that L^{A}TEX offers are not enough. An example is the following:

If a solution to an exercise goes across a page break, then I would like to have “(Continued on next page…)” at the bottom of the first page and “(Continued…)” at the top in the margin of the next page.

You cannot use L^{A}TEX’s mark mechanisms for this if you also want to use chapter and section information.

The code from figure 5 constitutes a package that gives you two extra marks that can be used in this situation^{11}. Here is a way to use this package:

\usepackage{extramarks}
...
\pagestyle{fancy}
\lhead{\firstxmark}
\rfoot{\lastxmark}
...
\extramarks{}{Continued on next page\ldots}
Some text that may or may not cross a page boundary...
\extramarks{Continued\ldots}{}

Note that the \extramarks command must be close to the text, i.e no empty lines (paragraph boundaries) should intervene. Otherwise the page may be broken at that boundary and the extramarks would come on the wrong page.

There are two new marks that can be used in the page layout with this package: If commands of the form \extramarks{m_{1}}{m_{2}} are given \firstxmark gives you the first m_{1} value and \lastxmark gives you the last m_{2} value of the current page. It also gives you the \firstleftmark and \lastrightmark commands that complement the standard L^{A}TEX marks.

To stress the point that marks are the correct way to do this, let me give you a “solution” that will not work^{12}:

\lhead{Continued}
\rfoot{Continued on next page\ldots}
Some text that may or may not cross a page boundary...
\lhead{}
\rfoot{}

You may be tempted to think that the first \lhead and \rfoot will be in effect when TEX breaks the page in the middle of the text, and the last ones when the page breaks after the text. This

11After I made this package I discovered a package secret.sty that does a similar thing to mark confidential paragraphs if they cross a page boundary. It does it, however, by changing the output routine.

12Actually there is another way but it requires two L^{A}TEX passes: you can put \label commands before and after the text and compare the \pagerefs.

% extramarks.sty
\def\@leftmark#1#2#3#4{#1}
\def\@rightmark#1#2#3#4{#2}
\def\markboth#1#2{{\def\protect{\noexpand\protect\noexpand}
     \let\label\relax \let\index\relax \let\glossary\relax
     \expandafter\@markboth\@themark{#1}{#2}
     \mark{\@themark}}\if@nobreak\ifvmode\nobreak\fi\fi}
\def\markright#1{{\def\protect{\noexpand\protect\noexpand}
     \let\label\relax \let\index\relax \let\glossary\relax
     \expandafter\@markright\@themark
     {#1}\mark{\@themark}}\if@nobreak\ifvmode\nobreak\fi\fi}
\def\@markright#1#2#3#4#5{\gdef\@themark{{#1}{#5}{#3}{#4}}}
\def\@markboth#1#2#3#4#5#6{\gdef\@themark{{#5}{#6}{#3}{#4}}}
\def\leftmark{\expandafter\@leftmark\botmark{}{}{}{}}
\def\rightmark{\expandafter\@rightmark\firstmark{}{}{}{}}
\def\firstleftmark{\expandafter\@leftmark\firstmark{}{}{}{}}
\def\lastrightmark{\expandafter\@rightmark\botmark{}{}{}{}}
\def\@themark{{}{}{}{}}
\def\extramarks#1#2{{\def\protect{\noexpand\protect\noexpand}
     \let\label\relax \let\index\relax \let\glossary\relax
     \expandafter\@markextra\@themark{#1}{#2}
     \mark{\@themark}}\if@nobreak\ifvmode\nobreak\fi\fi}
\def\@markextra#1#2#3#4#5#6{\gdef\@themark{{#1}{#2}{#5}{#6}}}
\def\firstxmark{\expandafter\@firstxmark\firstmark{}{}{}{}}
\def\topxmark{\expandafter\@firstxmark\topmark{}{}{}{}}
\def\lastxmark{\expandafter\@lastxmark\botmark{}{}{}{}}
\def\@firstxmark#1#2#3#4{#3}
\def\@lastxmark#1#2#3#4{#4}

Figure 5: Package for extra marks in L^{A}TEX

is not true as the whole paragraph (including the last definitions) will be processed before TEX considers the page break, so at the time of the page break the last definitions are effective, whether the page break occurs inside the text or outside of it. Putting a paragraph boundary between the text and the last definitions will not work either, because you don’t want the first definitions to be in effect when TEX decides to break the page exactly at this boundary. Actually the marks mechanism was invented to get rid of these kinds of problems.

In the above example the text “Continued” appears in the page header. It may be nicer to put it in the margin. This can be easily accomplished by positioning it at a fixed place relative to the page header. In plain TEX you would use a concoction of \hbox to 0pt, \vbox to 0pt, \hskip,\vskip, \hss and \vss but fortunately L^{A}TEX’s picture environment gives a much cleaner way to do this. In order not to disturb the normal header layout we put the text in a zero-sized picture. Generally this is the best way to position things on fixed places on the page. You can then also use the normal headings. See also section 20 for another example of this technique.

\lhead{\setlength{\unitlength}{\baselineskip}%
\begin{picture}(0,0)
  \put(-2,-3){\makebox(0,0)[r]{\firstxmark}}
  \end{picture}\leftmark}

This solution can of course also be used for the footer. Make sure you put the picture as the first thing in left-handside entries and last in right-handside ones.

Finally you may want to put “(Continued…)” in the text rather than in the header or the margin. Then you have to use the afterpage package. We also decide to make a separate environment for it.

\newenvironment{continued}{\par
  \extramarks{}{Continued on next page\ldots}
  \afterpage{\noindent\firstxmark\vspace{1ex}}
  }{\extramarks{(Continued\ldots)}{}\par}

It is a bit dangerous to use \firstxmark outside the page layout routine, but apparently with \afterpage this works. If you would need the information further on in the page you must remember the state of the marks in your own variable. You can set this in one of the fancyhdr fields. For example if you want to add something after the broken piece of text you can use the following:

\newcommand{\mysaved}{}
\newenvironment{continued}{\par
  \extramarks{}{Continued on next page\ldots}
  }{\extramarks{(Continued\ldots)}{}\par\vspace{1ex}\mysaved}
\lhead{\leftmark}
\chead{\ifthenelse{\equal{\lastxmark}{}}
  {\gdef\mysaved{}}
  {\gdef\mysaved{\noindent[Continued from previous page]}}}

If you want to include one of the marks or other varying information in the saved text, you must use \xdef rather than \gdef.

19 A movie

If you put at each page on the same place a picture that slightly changes from page to page you can get a movie-like effect by flipping through the pages. You can create such a movie easily with fancyhdr. For simplicity we assume that each picture is in a postscript (EPS) file called pic⟨n⟩.ps where ⟨n⟩ is the page number and that we use the graphics or graphicx package^{13}. To put the movie in the righthandside bottom corner the following will work:

\rfoot{\setlength{\unitlength}{1mm}
  \begin{picture}(0,0)
    \put(5,0){\includegraphics{pic\thepage.ps}}
  \end{picture}}

Notice that the \unitlength parameter should be set locally in the fancyhdr field in order to avoid unwanted interference with its value in the text.

20 Thumb-indexes

Some railroad guides and expensive bibles have so called thumb-indexes, i.e. there are marks on the sides of the pages that indicate where the chapters are. You can create these by printing black blobs in the margin of the pages. The vertical position should be determined by the chapter number or some other counter. As the position is independent of the contents of the page, we print these blobs as part of the header in a zero-sized picture as described in the previous section.

13If you use an older version of L^{A}TEX you could use the epsf or epsfig package.

Of course we have to take care of two-sided printing, and we may want to have an index page with all the blobs in the correct position. The solution requires some hand-tuning to get the blobs nicely spaced out vertically. For the application that I had there were 12 sections, so I made the blobs 18 mm apart, i.e. 9 mm blob separated by 9 mm whitespace. In order to avoid calculations they are set in a picture environment with the \unitlength set to 18 mm. Page numbers are set in the headers at the outer sides, and the blobs are attached to these. In this example the section numbers are used to position the blobs, but you can replace this with any numeric value. See figure 6 for the resulting overview page and figure 7 for the code.

1_{1 Introduction 2 The first year 3 Specialisation}

Figure 6: Thumb-index overview page

21 Float placement

Floats are page elements that float with respect to the rest of the document. Standard floats are tables and figures, but with the float package you can easily make new ones, like algorithms. Most of the time floats work satisfactory, but sometimes L^{A}TEX seems too stubborn to do what you want. This section describes how you can influence L^{A}TEX so that it will do most of the time what you want. There might, however be some pathological cases where it is impossible to convince L^{A}TEX to do things your way. In the following we will use figures as an example but everything applies to other floats as well.

The most encountered problems with floats are:

1. You want a float at a certain position in the text, but L^{A}TEX moves it, usually to the next page.

2. From a certain point, L^{A}TEX moves all your floats to the end of the document or the end of a chapter.

3. L^{A}TEX complains about “Too many floats”.

In the first two cases you must first check if you have given the correct “placement” parameter to you float, e.g. \begin{figure}[htp] specifies that your figure may be placed either: Here (i.e. in the text position where the command is given), on the Top of a page (which may be the page where you put the command), or on a separate Page of floats. You could also have specified “b” for Bottom of the page. The order of the letters is insignificant, you cannot force L^{A}TEX to try Bottom first and then Top by specifying [bt].

If L^{A}TEX doesn’t put the float at the place where you expected it, it is usually caused by the following:

\setlength{\unitlength}{18mm}
\newcommand{\blob}{\rule[-.2\unitlength]{2\unitlength}{.5\unitlength}}
\newcommand\rblob{\thepage
  \begin{picture}(0,0)
    \put(1,-\value{section}){\blob}
  \end{picture}}
\newcommand\lblob{%
  \begin{picture}(0,0)
    \put(-3,-\value{section}){\blob}
  \end{picture}%
  \thepage}
\pagestyle{fancy}
\cfoot{}
\newcounter{line}
\newcommand{\secname}[1]{\addtocounter{line}{1}%
  \put(1,-\value{line}){\blob}
  \put(-7.5,-\value{line}){\Large \arabic{line}}
  \put(-7,-\value{line}){\Large #1}}
\newcommand{\overview}{\thepage
  \begin{picture}(0,0)
    \secname{Introduction}
    \secname{The first year}
    \secname{Specialisation}
...etc...
  \end{picture}}
\begin{document}
\fancyhead[R]{\overview}\mbox{}\newpage % This produces the overview page
\fancyhead[R]{} % Front matter may follow here
\clearpage
\fancyhead[RE]{\rightmark}
\fancyhead[RO]{\rblob}
\fancyhead[LE]{\lblob}
\fancyhead[LO]{{\leftmark}
...

Figure 7: Thumb-index code 1. The float didn’t fit on the page. In this case it has to move to the next page or even further.

If you didn’t specify either [t] or [b] in the position parameter, L^{A}TEX must save it until it has enough for a page of floats. So don’t specify only [h]. If you want to give L^{A}TEX a chance to put the float on a page of floats, you must also specify “p”.

2. The placement would violate the constraints imposed by L^{A}TEX’s float placement parameters. This is one of the most occurring causes and it can easily be corrected by changing the parameters. Here is a list of them:

Counters – change with \setcounter

topnumber  max. number of floats at top of page  2
bottomnumber  max. number of floats at bottom of page  1
totalnumber  max. number of floats on a page  3

Other – change with \renewcommand

\topfraction  max fraction of page for floats at top  0.7
\bottomfraction  max fraction of page for floats at bottom  0.3
\textfraction  min fraction of page for text  0.2
\floatpagefraction min fraction of floatpage that should have floats 0.5

There are also some others for double column floats in two-column documents.

The values in the righthand column are the defaults for the standard L^{A}TEX classes. Other classes could use different defaults. As you see with the default values a float will not be put in the bottom of a page if its height is more than 30% of the page height. So if you specify [hb] for a float which is taller it has to move to a float page. But if it is less than 50% of the page height it will have to wait until some more floats are given before a float page can be filled to satisfy the \floatpagefraction parameter. If you have this kind of behaviour you can easily adapt the parameters, e.g. with:

\renewcommand{\textfraction}{0.05}
\renewcommand{\topfraction}{0.95}
\renewcommand{\bottomfraction}{0.95}
\renewcommand{\floatpagefraction}{0.35}
\setcounter{totalnumber}{5}

You may want to be careful not to make \floatpagefraction too small, otherwise you may get too many small floatpages.

You can force L^{A}TEX to ignore most of the parameters for one specific float occurrence by including an exclamation mark (!) in the placement parameters, e.g.

\begin{figure}[!htb]

Floats which contain a “t” in the position parameter could be placed before the place where they are referenced (but on the same page). This is normal behaviour for L^{A}TEX but some people just don’t like it. There are a number of ways to prevent this:

1. Of course deleting the “t” will help, but in general this is undesirable, as you may want the float to be placed at the top of the next page.

2. use the flafter package which causes floats never to be placed “backwards”.

3. use the command \suppressfloats[t]^{14}. This command will cause floats for the top position on this page to be moved to the next page. This can also be done with [b] or without parameter for all floats on this page.

14This command and the “!” placement parameter are not defined in L^{A}TEX2.09.

If in spite of all your attempts L^{A}TEX still moves your floats to the end of the document or the end of a chapter, you can insert a \clearpage command. This will start a new page and insert all pending floats before continueing. If it is undesirable to have a pagebreak you can use the afterpage package and the following command:

\afterpage{clearpage}

This will wait until the current page is finished and then flush all outstanding floats. In some pathological circumstances afterpage may give strange results, however.

Finally, if you want a float only at the place where you define it, without L^{A}TEX moving it whatsoever, you can use the float package and give the command:

\restylefloat{figure}

in the preamble. Now you will be able to specify [H] as the position parameter, which will mean “HERE and only HERE”. This may cause an unwanted page break however^{15}. If you want to avoid the unwanted pagebreak, i.e. let L^{A}TEX move the float only if it doesn’t fit on the page, the use the afterpage package with:

\afterpage{\clearpage \begin{figure}[H] ... \end{figure}}

Complaints from L^{A}TEX about “Too many floats” are usually caused by one of the above problems: floats not being able to be placed and L^{A}TEX collecting too many of them. The solutions given above, especially those with \clearpage in them will usually help. In some cases there really are too many floats, as L^{A}TEX has a limited number of “boxes” to store the floats. The package morefloats can be used to increase this number. If you need still more then you must edit a private copy of this file, but even then there will be some limit that you cannot pass. Then your only resort will be to change your document.

22 Multipage Floats

L^{A}TEX’s floats cannot be split across pages. Sometimes, however, you want to have a table or figure that doesn’t fit on one page. The easiest way is to split these into multiple table or figure environments, but this has a number of undesirable effects:

• Where do you split it? This is in general more difficult for tables than for figures.

• How do you keep them together?

• You don’t want more than one entry in the list of figures/tables.

Although these problems are not fully solvable in all cases, here are a couple of suggestions:

22.1 Tables

For tables longer than a page you can use the longtable package. This package defines a longtable environment that is a kind of amalgamation of table and tabular. It has approximately the same syntax as the tabular environment, but it adds some features of table, like captions. Longtables will be automatically split when they don’t fit on the page. And they will be entered in the list of tables when a caption is given. They will not float, however, and cannot be used inside a float environment. This could mean that another table environment, which was defined before the longtable, will float past it, and therefore the numbers may get out of order. Another problem could be that the longtable starts rather far down the page, which isn’t a pleasant sight. If you want the longtable to start at the top of the page, the best thing to do is to include it in an \afterpage command (using the afterpage package). As a longtable is by definition large, it is best to put it in a separate file, and \input it in the \afterpage command:

15There used to be a here.sty with the same effect, but this is incompatible with L^{A}TEX 2ε

\afterpage{\input{mytable}}
\afterpage{\clearpage\input{mytable}}

The last form has the additional advantage that most of the outstanding floats will be printed first.

22.2 Figures

There isn’t an equivalent longfigure solution, so for figures you will usually have to split it yourself. In general this is less of a problem. However, the problem you get now is how to keep them together, i.e. how to get the parts on subsequent pages, and how to get a single entry in the list of figures.

You will have to split the figure into pieces and put each part in a separate figure environment. To keep them together it is best to use only the [p] placement, so that they will be put on floatpages. As they are bigger than a page this is appropriate. The first part would then get a \caption, the subsequent parts would be used without a caption. If you want to add a caption-like text, enter it as normal text rather than a \caption, so that it will not be entered in the list of figures. It may also be desirable to issue a \clearpage first, just like we did for the longtable, and to encapsulate this in the \afterpage command. E.g.

\afterpage{\clearpage\input{myfigure}}

where myfigure.tex contains:

\begin{figure}[p]
\includegraphics{myfig1.eps}
\caption{This is a multipage figure}
\label{fig:xxx}
\end{figure}
\begin{figure}[p]
\includegraphics{myfig2.eps}
\begin{center}
Figure~\ref{fig:xxx} (continued)
\end{center}
\end{figure}

You have to make sure that the last part is big enough, otherwise L^{A}TEX could decide to postpone it until it has collected some more floats. This can be done either by making the figure big enough (e.g. by adding some \vspace), or by tweaking the \floatpagefraction parameter.

If you want your multipage figure to start at a lefthand-side (even-numbered) page you can use a test in the \afterpage command (using the ifthen package):

\afterpage{\clearpage
\ifthenelse{\isodd{\value{page}}{\afterpage{\input{myfigure}}} % odd page
    {\input{myfigure}}}} % even page

If there are too many floats on the skipped page, this may still fail to start your multipage figure on an even page, however.

23 Contact information

Piet van Oostrum Dept. of Computer Science Utrecht University P.O. Box 80.089 3508 TB Utrecht, The Netherlands Telephone: +31 30 2531806 Telefax: +31 30 2513791 E-mail: piet@cs.ruu.nl WWW: http://www.cs.ruu.nl/people/piet